# Optimal Inference of Asynchronous Boolean Networks

### v5

@@ Abstract
The network inference problem arises in biological research when one needs to quantitatively choose the best protein-interaction model for explaining a phenotype. The diverse nature of the data and nonlinear dynamics pose significant challenges in the search for the best methodology. In addition to balancing fit and model complexity, computational efficiency must be considered. In this paper, we present a novel approach that finds a solution that fits the observed dataset and otherwise a minimal number of unobserved datasets. We present algorithms for computing Boolean networks that optimally satisfy this criterion, and allow for asynchronicity network dynamics. Furthermore, we show that using our methodology a solution to the pseudo-time inference problem, which is pertinent to the analysis of single-cell data, can be intertwined with network inference. Results are described for real and simulated datasets.
@@
Boolean Network, Inference, Optimization, Time-series1 Introduction
@@
Research in molecular biology often aims to reveal a mechanistic understanding of underlying processes, in contrast to correlations between observations. In particular, one aims to describe cellular networks of interactions between proteins that generate a specific phenotype [1, 2, 3, 4]. Often, these networks display complex, nonlinear dynamics due to combinatorial interactions between regulators and feedback loops in the topology. [5]. A key aspect of the problem is the Curse of Dimensionality, which ensures that the number of networks and datasets grows exponentially with the number of involved proteins, making inference susceptible to overfitting.
@@
Gene regulatory networks, which control the expression of genes in the cell, are an important type of biological networks that are studied using transcriptomic technologies. These technologies measure the level (expression) of all the genes in a sample and, more recently, in single cells [6], allowing the detection of regulatory relationships between them. When the measurement is at the sample level, we refer to the data as RNA-Sequencing data (RNA-Seq), and when it is at the single-cell level as single-cell RNA-Seq (scRNA-Seq).
@@
Numerous, fundamentally different inference methodologies have been proposed. For example, Keyl et al. used an explainable AI approach and layer-wise relevance propagation in order to select gene regulators that are ranked as important predictors in a neural network model [7]. SCODE is a method for inferring ordinary differential equations from scRNA-Seq data, such that the regulatory relationships can be derived from the equations [8]. CEFCON uses a graph neural network with an attention mechanism for constructing the regulatory network from prior network knowledge, scRNA-Seq data and trajectory information [9]. Chen et al. used a genetic algorithm to reconstructed a Probabilistic Boolean Network from scRNA-Seq data, a model that accommodates probabilistic transitions in a Boolean network model [10]. However, none of the existing methodologies satisfy all of the following criteria: define an objective criterion for comparing any pair of networks, find optimal solution with respect to this inference criterion, regularize any solution as part of the inference criterion and process both steady-state and time-series data. Presenting such a methodology is the goal of this paper.
In fitting a model that can be described using discrete quantities, and in the absence of dataset noise, the best solution is the one that uses the shortest encoding, since with the strictest limit on encoding length one can fit the smallest number of unobserved datasets. This minimal length encoding is known as the Kolmogorov Complexity [11]. In the case of biological datasets, noise is practically always present, but the same principle can be used for inference in the presence of noise. We will show that this leads to an inference methodology that satisfies the requirements described above, using Boolean networks as the modeling language [12, 13, 14].
@@
Another challenge that arises specifically when working with scRNA-Seq is that individual measurements originate from a time-series but a temporal order is not provided with the dataset. In other words, we obtain "snapshots" of the cell at different states, that may have occurred consecutively in time and thus describe a network trajectory, but we do not know at which point in the trajectory each state, or snapshot, occurred. Here, too, many efforts have been devoted to the development of inference methods such that the cell order, known as pseudo-time, could be derived. To name a few examples, Riba et al. developed DeepCycle, a method that predicts position relative to the cell cycle based on the relationship between spliced and unspliced transcripts. DeepCycle trains an autoencoder that encodes the relationship between spliced and unspliced scRNA-Seq reads as the transcriptional phase of the cell, a latent variable that maps a cell into a position on the periodic cell cycle trajectory [15]. Liu et al. introduced reCAT, a method for cell-cycle stage inference that finds a cycle that minimizes transcriptional differences between adjacent cells [16]. The CONFESS method utilizes image information to reconstruct single-cell dynamics [17]. PAGA creates a map connecting cell groups based on the estimation of the connectivity of manifold partitions [18]. GraphDDP combines clusters into differentiation trajectories using a forced-based graph layout approach [19]. We argue that the pseudo-time inference problem is closely related to the network inference problem and that a network model should be utilized to solve it. We will demonstrate how this can be done in the context of the novel methodology suggested here.
@@ 2.1 Network Inference
We will also assume that a set of plausible regulatory interactions has been determined for each gene. This can be achieved using other experimental technologies like ChIP-Seq [20], or by keeping regulators whose levels correlate with their target’s.
From this initial set, the modeler want to choose the optimal subset, including the logic tables by which the regulators determine the state of the target. For mapping biological measurements to Boolean values we will use existing methods [21, 22, 23, 24], and in this section we assume that all gene values are already Boolean.
Were the data noiseless, we would seek a solution whose binary encoding is as short as possible, since the number of networks that can be described by such an encoding is minimal, and hence the solution will fit a minimal amount of unobserved data [11]. Every bit doubles the number of networks that the encoding can describe. In practice, datasets are always affected by noise. We therefore need a criterion for combining noise bits and network bits such that the solution will match a minimal number of unobserved datasets. Karlebach and Robinson [25] showed, in the context of synchronous dynamics, that minimizing the sum of network encoding bits and noise bits is equivalent to minimizing the number of unobserved datasets that can be matched. A generalization to asynchronous dynamics will be discussed shortly.
@@
Now if we set the  variable to 0 instead, set the suffix of the trajectory from time  to fit exactly the trajectory of the model that was fit to  from time  and afterwards, then  must also be optimal for this new trajectory . If not, and there is another better fit model, then when flipping the bits back, it will still be better than  on the original dataset, whether  uses the  variable or not. Therefore, like the  variables, every  variable is equivalent to one bit in the encoding of the network. The value of the objective function is a sum of the weights of variables that are set to 1 in the solution.
Powerful solvers like Gurobi [26] have dramatically improved our ability to solve 0/1 Integer Programming problems. Custom heuristics can be integrated with the solver to improve performance. We now describe such heuristics.
Perhaps the simplest heuristic for a trajectory is to perform a single pass over the data, state by state starting from the first state, and to record every input-output pair observed as long as it does not conflict with pairs observed before it. When a conflict occurs, the value of the target gene is flipped to match the output that was previously observed. A more sophisticated approach was suggested by Karlebach and Robinson [25], and can be applied to an expression data set composed of either steady states or equal-length trajectories:
@@ 2.2 Pseudo-time
Psuedo-time assignment, also known as trajectory inference, is the ordering of individual states, produced by scRNA-Seq experiments, into trajectories and steady states. The method described in the previous section can be adopted for this purpose, in a procedure that resembles Expectation-Maximization. Initially, the reconstructed network is set to the model that contains all the candidate edges, with some initial logic. For example, if edge signs are available or estimated from correlations, a rule like "inhibitors win over activators’ can be used. Next, using the initial network, network trajectories from the states corresponding to the observed Boolean states of the cells are generated: each trajectory is extended until an attractor is reached, and also includes all attractor states. Each observed Boolean cell state is then mapped to its closest state in the trajectories that were obtained. Contiguous sequences of states are assigned to the corresponding trajectories time points, where trajectories of length 1 are set as steady states. Using the assignment of pseudo-time, a new network is reconstructed from the data, and a set of de-noised cell states is obtained as part of this reconstruction. Each of these can be used as an initial state for the inferred network in order to generate trajectories, map the data to them, and obtain a new pseudo-time. Using the new pseudo-time, a network is fitted to the data. The process repeats itself until the value of the solution stops improving. In the EM analogy, the latent variables are the pseudo-times assigned to cell states, and the network edges and logic are the model parameters. For ease of reference we will refer to this procedure as TICO (Timeless Inference of Cell Ordering).
@@ 3.1 Synchronized Cells
To test our method on real experimental data, we obtained the microarray dataset GSE49650 of synchronized yeast cells from the Gene Expression Omnibus. Preprocessing of the .CEL files was done using the rma function in the Bioconductor package affy, using default argument values. Additionally, every trajectory (time-series) of every gene was smoothed using the R functions smooth.spline with smooth parameter 0.5 and then approxfun (in the stats package) with default argument values [27]. The x-coordinate values for the smoothing were the times at which the measurements were taken in minutes, and the y-coordinate values were the array intensities. We used the BASCA method as implemented in the R package Binarize [21], with default arguments, for mapping from continuous to Boolean values. Every trajectory was binarized separately.
We used the yeast cell cycle model from Cho et al. [28], with the edges as the candidate regulatory connections. Complexes were modeled using the expression of one of their genes. After fitting the model, the percentage of mismatches was 16.3 %.
In Cho et al., edges were associated with activation or repression activity. Compared to the inferred logic tables, all the edges agreed on the sign with Cho et al. This result is detailed in Table 1. The inferred logic provided information about combinatorial regulations, i.e. what is the joint effect of multiple regulators. This is illustrated in Figure 1 for two of the targets. Figure 2 shows an inferred trajectory for one of the time-series, and the original continuous levels for two of the genes. Interestingly, genes can be separated into those that are active at early or late cell cycle stages.
It should be noted that not every candidate interaction was inferred. The reason for this is that either the dataset size is too small for obtaining reliable predictions, or that the interactions occurring in the dataset do not capture all the candidate interactions. Both the synchronous and asynchronous dynamics of the inferred model lead to the same steady state, suggesting a robust design. Perturbing a single gene in the steady state generates an oscillation that settles back to the steady state.
@@
Next, we applied the pseudo-time inference algorithm (TICO) to the HSC network of Bonzanni et al. [29], using GEO dataset GSE75478 [30], which profiles human hematopoeitic stem cells in early differentiation from the bone marrow of two individuals.
The raw data was preprocessed using the R package scuttle, followed by library-size normalization and log transformation. Binarized values were obtained using the BASCA method [21] from the R package Binarize, with default arguments. As an initial network for the pseudo-time inference algorithm, we used the logic published in Bonzanni et al. For validation, we used the CD38 measurements for individual cells, as CD38 is a marker for differentiation and therefore expected to increase along differentiation trajectories.
The binarized cell states were mapped to their closest counterparts in the network trajectories for the initial and inferred networks, and Pearson correlation and the corresponding correlation test p-values were computed between CD38 measurements in consecutive states. As shown in figure 3, correlation values from the inferred model were higher, with fold changes of 1.7 and 2.6 for individuals 1 and 2, respectively. The corresponding p-values were also lower, and significant for the inferred networks -  for individual 1 and  for individual 2. The  ratio of the p-values obtained for the inferred and the initial networks are displayed in figure 3.
@@ 3.3 Simulation
To test TICO using a known ground-truth, we generated 20 simulated datasets from networks the size of Bonzanni et al. Two random true regulators and one random false regulator, and an interaction between the false regulator and true ones such that it has maximal agreement with the data, were added for each gene. In order to generate count data that is similar to scRNA-Seq experiments, we binarized the real data from GEO dataset GSE75478 to obtain active and inactive labels, and fitted each gene with a zero-inflated negative binomial model. Network dynamics were generated using the R package BoolNet [31], with some asynchronous updates, and Boolean values were mapped to counts by generating random values from the fitted models of randomly chosen genes in GSE75478.
For comparison, we used SCORPIUS [32] to infer pseudo-time and the bestfit method from the Boolnet [31] package to reconstruct the networks. Since the pseudo-time generated by SCORPIUS does not assign the cells to individual trajectories, we provided SCORPIUS with the trajectory labels - only the order of cells/states within trajectories needed to be inferred. Additionally, SCORPIUS does not provide directionality in time. Therefore, for SCORPIUS predictions we chose the directionality that agrees best with the true trajectory. These choice provide a significant boost to SCORPIUS predictions - for example, trajectories of length 2 will always be inferred correctly. Nevertheless, the predictions obtained by TICO were significantly better. Figure 4 shows the true positives (y-axis) and false positives (x-axis) for different datasets obtained by the two methods(blue- TICO, red - SCORPIUS+bestfit). The green line is the ratio of TP and FP expected by chance. As can be seen in the figure, the blue dots are significantly removed from the green line, whereas the red ones are spread along its margin. When computing the probability of randomly drawing the observed number of TPs or more, the upper tail of the hypergeometric distribution is less than 0.05 for 15 of reconstructions obtained by TICO, and only for one reconstruction obtained by SCORPIUS+bestfit.

### v6

@@ Abstract
The network inference problem arises in biological research when one needs to quantitatively choose the best protein-interaction model for explaining a phenotype. The diverse nature of the data and nonlinear dynamics pose significant challenges in the search for the best methodology. In addition to balancing fit and model size, computational efficiency must be considered. Importantly, underlying the measurements, which are affected by experimental noise, there is a complex computational mechanism that is inherently hard to identify. To address these difficulties, we present a novel approach that uses algorithmic complexity to infer a Boolean network model from experimental data. We present an algorithm that is optimal within this framework and allows for asynchronicity network dynamics. Furthermore, we show that using our methodology a solution to the pseudo-time inference problem, which is pertinent to the analysis of single-cell data, can be intertwined with network inference. Results are described for real and simulated datasets.
@@
Boolean Network, Inference, Optimization, Time-series,Gene Regulatory Network1 Introduction
@@
Research in molecular biology often aims to reveal a mechanistic understanding of underlying processes, in contrast to correlations between observations. In particular, one aims to describe cellular networks of interactions between proteins and other molecules that generate a specific phenotype [Karlebach2008, Arenas2015, Hackett2016, Eichenberger2004]. Often, these networks display complex, nonlinear dynamics due to combinatorial interactions between regulators and feedback loops in the topology. [Kauffman1969]. A key aspect of the problem is the Curse of Dimensionality, which ensures that the number of networks and datasets grows exponentially with the number of involved proteins, making inference susceptible to overfitting.
@@
Gene regulatory networks, which control the expression of genes in the cell, are an important type of biological networks that are studied using transcriptomic technologies. These technologies measure the level (expression) of all the genes in a sample and, more recently, in single cells [Andrews2020], allowing the detection of regulatory relationships between them. When the measurement is at the sample level, we refer to the data as RNA-Sequencing data (RNA-Seq), and when it is at the single-cell level as single-cell RNA-Seq (scRNA-Seq).
@@
Numerous, fundamentally different inference methodologies have been proposed. For example, Keyl et al. used an explainable AI approach and layer-wise relevance propagation in order to select gene regulators that are ranked as important predictors in a neural network model [Keyl2023]. SCODE is a method for inferring ordinary differential equations from scRNA-Seq data, such that the regulatory relationships can be derived from the equations [Matsumoto2017]. CEFCON uses a graph neural network with an attention mechanism for constructing the regulatory network from prior network knowledge, scRNA-Seq data and trajectory information [Wang2023]. Chen et al. used a genetic algorithm to reconstructed a Probabilistic Boolean Network from scRNA-Seq data, a model that accommodates probabilistic transitions in a Boolean network model [Chen2014]. However, none of the existing methodologies satisfy all of the following criteria: define an objective criterion for comparing any pair of networks, find optimal solution with respect to this inference criterion, regularize any solution as part of the inference criterion and process both steady-state and time-series data. Presenting such a methodology is the goal of this paper.
We will show that using a computational definition of randomness leads to an inference methodology that satisfies the requirements described above, using Boolean networks as the modeling language [Fox2024, Samaga2009, Schwab2017].
@@
Another challenge that arises specifically when working with scRNA-Seq is that individual measurements originate from a time-series but a temporal order is not provided with the dataset. In other words, we obtain "snapshots" of the cell at different states, that may have occurred consecutively in time and thus describe a network trajectory, but we do not know at which point in the trajectory each state, or snapshot, occurred. Here, too, many efforts have been devoted to the development of inference methods such that the cell order, known as pseudo-time, could be derived. To name a few examples, Riba et al. developed DeepCycle, a method that predicts position relative to the cell cycle based on the relationship between spliced and unspliced transcripts. DeepCycle trains an autoencoder that encodes the relationship between spliced and unspliced scRNA-Seq reads as the transcriptional phase of the cell, a latent variable that maps a cell into a position on the periodic cell cycle trajectory [Riba2022]. Liu et al. introduced reCAT, a method for cell-cycle stage inference that finds a cycle that minimizes transcriptional differences between adjacent cells [Liu2017]. The CONFESS method utilizes image information to reconstruct single-cell dynamics [Motakis2018]. PAGA creates a map connecting cell groups based on the estimation of the connectivity of manifold partitions [Wolf2017]. GraphDDP combines clusters into differentiation trajectories using a forced-based graph layout approach [Costa2018]. We argue that the pseudo-time inference problem is closely related to the network inference problem and that a network model should be utilized to solve it. We will demonstrate how this can be done in the context of the novel methodology suggested here.
@@ 2.1 Network Inference
We will also assume that a set of plausible regulatory interactions has been determined for each gene. This can be achieved using other experimental technologies like ChIP-Seq [Mathelier2013], or by keeping regulators whose levels correlate with their target’s.
From this initial set, the modeler want to choose the optimal subset, including the logic tables by which the regulators determine the state of the target. For mapping biological measurements to Boolean values we will use existing methods [Hofensitz2012, Glaz2001, Shmulevich2002, Berestovsky2013], and in this section we assume that all gene values are already Boolean.
The methodology that we present is based on the concept of Kolmogorov complexity [Kolmogorov1968]. Randomness is not defined using probabilities, but rather as datasets that cannot be generated by a Boolean network with a significantly shorter encoding. Since the number of Boolean networks with a significantly shorter encoding is exponentially smaller, only datasets that are not random will have a model that can generate them. This leads to an optimization criterion: find a network that can be encoded with the smallest number of bits, that matches as many bits as possible in the dataset, where all bits have the same weights.
One has to note that datasets may only contain any trajectory or steady state once. The question that we ask is whether these could be generated by a concise Boolean network, and thus multiple occurrences of the exact output do not alter the randomness or non-randomness of the dataset, just as running a computer program again and getting the same output does not change our conclusions about the nature of the program.
Another potential source of randomness can be the computation itself, i.e. when the network displays non-determinism in shifting between states. Every non-deterministic step branches between two possible states, and so intuitively doubles the number of states generated by the network. The number of such steps can then be added to the sum that we wish to minimize.
@@
Now if we set the variable to 0 instead, set the suffix of the trajectory from time to fit exactly the trajectory of the model that was fit to from time and afterwards, then must also be optimal for this new trajectory . If not, and there is another better fit model, then when flipping the bits back, it will still be better than on the original dataset, whether uses the variable or not. Therefore, like the variables, every variable is equivalent to one bit in the encoding of the network. The value of the objective function is a sum of the weights of variables that are set to 1 in the solution, i.e.:
where W is a function that maps between a V variable and the weight assigned to it. This concludes the 0/1 Integer Programming formulation.
Powerful solvers like Gurobi [gurobi] have dramatically improved our ability to solve 0/1 Integer Programming problems. Custom heuristics can be integrated with the solver to improve performance. We now describe such heuristics.
Perhaps the simplest heuristic for a trajectory is to perform a single pass over the data, state by state starting from the first state, and to record every input-output pair observed as long as it does not conflict with pairs observed before it. When a conflict occurs, the value of the target gene is flipped to match the output that was previously observed. A more sophisticated approach was suggested by Karlebach and Robinson [Karlebach2023], and can be applied to an expression data set composed of either steady states or equal-length trajectories:
@@
2.2 Computational Complexity
For a Boolean network with N genes, where each gene has K potential regulators, there are possible structures, and for each such structure on average about logics. For N=10 genes and K=3, this gives possible networks. Since this makes brute-force searches impractical, in this section we discuss the relationship between input size and running time. In [Karlebach2012] the Vertex Cover problem is reduced to the network inference problem by mapping each vertex to a regulator, and each edge that it covers to a target. The network structure is assumed to be fixed, and the noise bits (represented by the variables) are minimized. Since the resulting dataset consists of steady states, it applies to both synchronous and asynchronous networks. If we don’t assume that the network structure is fixed, we can simply duplicate each target and its incoming edges multiple times, such that the dataset will be large enough to derive the network topology. Hence, the problem described in this paper is NP-Complete. In practice, higher levels of noise will result in longer running times, whereas for noiseless datasets the minimal network can be found relatively fast. Interestingly however, the dataset constructed by the reduction in [Karlebach2012] can be modified not to contain noise. That is, once the network structure has been selected, none of the measured gene values needs to be changed. This can be done by changing every regulator value in one of the two experiments. In this modified reduction, the vertex cover is the set of nodes that correspond to regulators selected for the network structure. Hence, assuming , in the worst case the structure inference alone will require exponential time. Considering the fast that the computation of the optimal network is performed in-silico, we argue here that the investment of computational resources is worthwhile. Once a network has been inferred, many insights that could have taken a significant amount of wet lab work could be derived. The Simulation section describes the gains in accuracy that are obtained by solving this problem. It is also important to notice that when using an ILP solver, the good part of the computation is devoted to proving the optimality of a solution after the optimal solution has been found, so often an optimal solution can be found even if a time limit is imposed.
@@ 2.3 Pseudo-time
Psuedo-time assignment, also known as trajectory inference, is the ordering of individual states, produced by scRNA-Seq experiments, into trajectories and steady states. The method described in the previous section can be adopted for this purpose, in a procedure that analogous to Expectation-Maximization. Initially, the reconstructed network is set to the model that contains all the candidate edges, with some initial logic. For example, if edge signs are available or estimated from correlations, a rule like "inhibitors win over activators’ can be used. Next, using the initial network, network trajectories from the states corresponding to the observed Boolean states of the cells are generated: each trajectory is extended until an attractor is reached, and also includes all attractor states. Each observed Boolean cell state is then mapped to its closest state in the trajectories that were obtained. Contiguous sequences of states are assigned to the corresponding trajectories time points, where trajectories of length 1 are set as steady states. Using the assignment of pseudo-time, a new network is reconstructed from the data, and a set of de-noised cell states is obtained as part of this reconstruction. Each of these can be used as an initial state for the inferred network in order to generate trajectories, map the data to them, and obtain a new pseudo-time. Using the new pseudo-time, a network is fitted to the data. The process repeats itself until the value of the solution stops improving. In the EM analogy, the latent variables are the pseudo-times assigned to cell states, and the network edges and logic are the model parameters. To understand the procedure better under the framework of algorithmic complexity, recall that we start with a fixed network structure and order the states to fit its trajectories. For simplicity consider a single trajectory. As the number of states M increases, the possible ways to order the states increases but constitutes an increasingly smaller fraction of all possible datasets. Hence, for an arbitrary single network the oder of the states would form a random sequence of Boolean vectors. For a network whose dynamics are related to those of the real network, we will obtain non-random sequences of consecutive states, such that if we account for noise we can infer the real network more closely. By repeating the process, we converge to the real network. Convergence depends on starting "close-enough" to the real network. For ease of reference we will refer to this procedure as TICO (Timeless Inference of Cell Ordering).
@@ 3.1 Synchronized Cells
To test our method on real experimental data, we obtained the microarray dataset GSE49650 of synchronized yeast cells from the Gene Expression Omnibus. Preprocessing of the .CEL files was done using the rma function in the Bioconductor package affy, using default argument values. Additionally, every trajectory (time-series) of every gene was smoothed using the R functions smooth.spline with smooth parameter 0.5 and then approxfun (in the stats package) with default argument values [Becker1988]. The x-coordinate values for the smoothing were the times at which the measurements were taken in minutes, and the y-coordinate values were the array intensities. We used the BASCA method as implemented in the R package Binarize [Hofensitz2012], with default arguments, for mapping from continuous to Boolean values. Every trajectory was binarized separately.
We used the yeast cell cycle model from Cho et al. [Cho], with the edges as the candidate regulatory connections. Complexes were modeled using the expression of one of their genes. After fitting the model, the percentage of mismatches was 16.3 %.
In Cho et al., edges were associated with activation or repression activity. Compared to the inferred logic tables, all the edges agreed on the sign with Cho et al. This result is detailed in Table 1. The inferred logic provided information about combinatorial regulations, i.e. what is the joint effect of multiple regulators. This is illustrated in Figure 1 for two of the targets. Figure 2 shows an inferred trajectory for one of the time-series, and the original continuous levels for two of the genes. Interestingly, genes can be separated into those that are active at early or late cell cycle stages. More specifically, CDH1, MCM1, SWI4 and CLN3 are active at early time points, in accordance with their association with the G1 phase[Cho]. The other genes are active at later time points, in accordance with their association with S/G2/M.
It should be noted that not every candidate interaction was inferred. The reason for this is that either the dataset size is too small for obtaining reliable predictions, or that the interactions occurring in the dataset do not capture all the candidate interactions. Perturbing a single gene in the steady state generates an oscillation that settles back to the steady state.
@@
Next, we applied the pseudo-time inference algorithm (TICO) to the HSC network of Bonzanni et al. [Bonzanni2013], using GEO dataset GSE75478 [Velten2017], which profiles human hematopoeitic stem cells in early differentiation from the bone marrow of two individuals.
The raw data was preprocessed using the R package scuttle, followed by library-size normalization and log transformation. Binarized values were obtained using the BASCA method [Hofensitz2012] from the R package Binarize, with default arguments. As an initial network for the pseudo-time inference algorithm, we used the logic published in Bonzanni et al. For validation, we used the CD38 measurements for individual cells, as CD38 is a marker for differentiation and therefore expected to increase along differentiation trajectories.
The binarized cell states were mapped to their closest counterparts in the network trajectories for the initial and inferred networks, and Pearson correlation and the corresponding correlation test p-values were computed between CD38 measurements in consecutive states, and thus correctly describes the phenotype. As shown in figure 3, correlation values from the inferred model were higher, with fold changes of 1.7 and 2.6 for individuals 1 and 2, respectively. A positive correlation corresponds to increse in CD38 levels as cells progress in their trajectories. The corresponding p-values were also lower, and significant for the inferred networks - for individual 1 and for individual 2. The ratio of the p-values obtained for the inferred and the initial networks are displayed in figure 3.
@@ 3.3 Simulation
In order to evaluate the accuracy of our inference algorithm, we generated 20 random networks using the R package BoolNet [BoolNet], and generated both steady-states and trajectories to test the algorithm and other methods that use steady states or trajectories. The trajectories include asynchronous updates and noise, where as steady states do not, as by definition they cannot include asynchronous updates. For tools that require continuous data, Boolean 0 and 1 are mapped to normally distributed continuous values -1 and 1, respectively, with standard deviation 0.25. The length of each trajectory depends on the network - the simulation includes all trajectory states until a state reoccurs.
The other tools that we used are CLR [Faith2007], ARACNE [Margolin2006], GENIE3 [HuynhThu2010] , BESTFIT [Lhdesmki2003] and GENENET [OpgenRhein2007]. As can be seen in figure4, our approach (denoted as MEDSI) is significantly more accurate than other methods. This boost in accuracy is likely due to the fact that our approach takes into account the computation performed by the networks, which allows a much more efficient separation between data and noise.